OpenStudy (anonymous):

A line has a slope of 9/4 and a y-intercept of −14. Which equations represent the line? Choose all answers that are correct. A.-9x+4y=-56 B.9x+4y=56 C.y=9/4x-14 D.y=9/4x+14

OpenStudy (anonymous):

The slope represents x and y intercept represents y, therefore the slope is y=9/4x -14

OpenStudy (anonymous):

are there any more?

OpenStudy (anonymous):

A.-9x+4y=-56 is another equation

OpenStudy (anonymous):

If you put -9x+4y=-56 in a slop intercept form it would be y=9/4x -14

OpenStudy (anonymous):